# Impedance-Engineered Josephson Parametric Amplifier with Single-Step Lithography

### v1

@@
We present the experimental demonstration of an impedance-engineered Josephson parametric amplifier (IEJPA) fabricated in a single-step lithography process. Impedance engineering is implemented using a lumped-element series LC circuit. We use a simpler lithography process where the entire device – impedance transformer and JPA – are patterned in a single electron beam lithography step, followed by a double-angle Dolan bridge technique for Al-AlOx-Al deposition. We observe nearly quantum-limited amplification with 18 dB gain over a wide 400400400\,400MHz bandwidth centered around 5.35.35.3\,5.3GHz, and a saturation power of $-114\,$dBm. To accurately explain our experimental results, we extend existing theories for impedance-engineered JPAs to incorporate the full sine nonlinearity of both the JPA and the transformer. Our work shows a path to simpler realization of broadband JPAs and provides a theoretical foundation for a novel regime of JPA operation.
@@
Josephson Parametric Amplifiers (JPAs) Yurke et al. (1989) have been a mainstay of superconducting circuit QED systems Vijay, Slichter, and Siddiqi (2011); Hatridge et al. (2013); Lin et al. (2013); Roy and Devoret (2016) due to their ability to provide quantum-limited amplificationCaves (1982), facilitate single-shot readoutKrantz et al. (2016), generate squeezed states of lightMallet et al. (2011); Yurke et al. (1988), and to generate two-mode squeezed states through three-wave and four-wave mixing mechanismsEichler et al. (2011). The phase-insensitive amplification mode of these amplifiers is usually characterized by their gain-bandwidth product defined by the maximum gain times the full width at half maximum of the amplifier. Conventional JPAs are usually narrow-band amplifiers with low gain-bandwidth products (202020\,20dB gain over $20-50\,$MHz) Castellanos-Beltran and Lehnert (2007); Yamamoto et al. (2008); Mutus et al. (2013); Zhou et al. (2014), which limit their utility to amplifying a signal composed of a single mode/tone (frequency component). On the other hand, amplifiers with higher gain-bandwidth products can be used to amplify a range of frequencies over a wider bandwidth. This finds applications in multiplexing qubit readout by allowing the amplification of multiple readout tones within the broad bandwidth of the JPA Vijay (2020).
@@
In order to increase the bandwidth without compromising on the gain, and thereby boost the gain-bandwidth product, Travelling Wave Parametric Amplifiers (TWPAs) CULLEN (1958); Feldman, Parrish, and Chiao (1975); Wahlsten, Rudner, and Claeson (1977); Ho Eom et al. (2012); Yaakobi et al. (2013); Zorin (2019) were developed. TWPAs are transmission lines with a distributed nonlinear inductance that typically provide amplification of around 202020\,20dB over a 444\,4GHz bandwidth. However, they involve complex fabrication techniques that are required to create thousands of identical unit cells with Josephson junctions (JJs) and capacitors Macklin et al. (2015); White et al. (2015).
@@
In this Letter, we demonstrate an impedance-engineered JPA optimized for four-wave mixing-based amplification that is fabricated using a single-step lithography technique. The simplicity of our process, in turn, reduces the cost and time needed to fabricate the device. It also reduces the errors in fabrication and thereby increases the reliability and yield of the devices. The device used in this work has a simple design made up entirely of coplanar elements. Operating the device in reflection mode, we demonstrate a high gain over a wide bandwidth with almost quantum-limited noise performance, comparable to state-of-the-art devices Roy et al. (2015); Grebel et al. (2021); Lu et al. (2022). Furthermore, to explain our results, we develop a theoretical model to compute the expected gain of an impedance-engineered JPA accounting for the full non-linearity of both the JPA and the impedance transformer. Our model is a substantial extension of previous theoretical work where only the lowest order non-linearity of the JPA is considered and the transformer is assumed to be made of a linear inductor.
@@
Figure 1 shows the false color micrograph of the device and the zoomed-in images of an array of JJs of the transformer and the JPA SQUID. The JPA is formed by a dc-SQUID (nominal area of $138\,\mu$m2 and unbiased critical current of $1.8\,\mu$A) shunting an interdigital capacitor (IDC) with a simulated capacitance $C_{J}\approx 1.6\,$pF to ground. The impedance transformer, comprising an array of three nominally identical Josephson junctions, each with a critical current of 585585585\,585nA, connected in series with another IDC (with a simulated capacitance of $C_{T}\approx 420\,$fF), was placed in series between the transmission line and JPA, forming a galvanic connection with both structures. The JPA and transformer are shown in the false color micrograph in yellow and blue color respectively. To design the device layout, we used Qiskit Metal and we performed eigenmode simulations of the JPA and the impedance transformer using ANSYS HFSS. For estimating the capacitance of the JPA and the impedance transformer, we used ANSYS HFSS along with COMSOL MULTIPHYSICS. Using these simulations, in addition to the capacitances quoted above, we estimate the geometric inductance for the JPA and the transformer to be 0.120.120.12\,0.12nH and 0.200.200.20\,0.20nH, respectively.
The device was fabricated on an intrinsic silicon substrate using an aluminum evaporative deposition process. First, we did a preliminary cleaning of the substrate using “piranha” (a $3:1$ volume-ratio mixture of $\text{H}_{2}\text{SO}_{4}$ and $\text{H}_{2}\text{O}_{2}$, for 10 minutes) followed by dilute hydrofluoric acid(HF) – a $50:1$ volume-ratio mixture of DI water and HF, for 30 seconds. Next, a resist bi-layer comprising a 500500500\,500nm thick bottom layer of methacrylic acid (MAA) in methyl methacrylate (MMA), diluted in ethyl lactate (EL) at a 11% concentration, and a 200200200\,200nm thick top layer of 4% PMMA 950K in Anisole was spin-coated on the substrate. We then carried out e-beam lithography in a Raith e-Line at 202020\,20kV acceleration voltage. We pattern both the impedance transformer and the JPA in this step. The pattern development was done using the MIBK (Methyl Isobutyl Ketone) developer. Subsequently, we used the Dolan-bridge technique for forming the Josephson junctions through the double-angle (shadow) evaporation of aluminum. The double-angle deposition was done at ±$\pm 45^{\circ}$ with the first and second layers of Al having thicknesses 202020\,20nm and 404040\,40nm, respectively. All the Josephson junctions in the circuit – both in the JPA SQUID loop and the impedance transformer – were formed in a single in-situ oxidation step between the double-angle Al depositions. The oxidation was done at 700 mTorr pressure for 20 minutes.
The device was packaged inside a copper sample box using gold wire bonds. The sample box was then mounted on the mixing chamber plate of a dilution refrigerator at 151515\,15mK. To tune the frequency of the JPA, we applied a DC magnetic flux using an external coil made of superconducting wire. The sample box containing the JPA device was mounted at the center of the external coil such that the plane of the SQUID of the JPA was perpendicular to the direction of the magnetic field of the coil. The resonance frequency $\Omega_{J}$ of the JPA could then be varied using the external magnetic flux $\Phi_{\text{DC}}$. To further prevent noise due to external magnetic fields, the sample box, along with the magnetic coil, was placed inside a superconducting aluminum shield mounted at the mixing chamber.
@@
$n_{\text{add}}=T_{\text{HEMT}}\Bigl{[}\frac{\text{SNR}_{\text{off}}}{\text{SNR}_{\text{on}}}-\frac{L}{G}\Bigr{]}\frac{k_{b}}{\hbar\omega},$ (1)
where $n_{\text{add}},T_{\text{HEMT}}$ are the added noise photons by JPA and linearised HEMT noise temperature respectively, $\text{SNR}_{\text{off}},\text{SNR}_{\text{on}}$ are the measured SNR when the pump is off and on respectively, $G$ is the gain of the JPA, and $L$ is the loss between the JPA and the HEMT. In our analysis, we assume that the $T_{\text{HEMT}}$ is equal to the nominal noise temperature in the datasheet of 3.63.63.6\,3.6K, and we estimate the loss between the IEJPA and the HEMT to be about 1.81.81.8\,1.8dB, mostly attributed to insertion loss for the circulators and the cable loss. The plot of estimated $n_{\text{add}}$ is shown by the orange curve in the middle panel of Fig. 2. We note that for a more precise characterization of the added noise, we need to use a shot-noise tunnel junction (SNTJ) noise thermometerSimoen et al. (2015), which was not available during this work. The green dotted line in Fig. 2(b) corresponds to the standard quantum limit of 0.50.50.5\,0.5 photons of noise added.
Figura 2: Measured (a) gain, (b) added noise, and (c) 1dB compression point for the IEJPA operated at a pump frequency and pump power of 5.3425.3425.342\,5.342GHz and $-88\,$dBm respectively. In the plot (b), the solid line represents the estimated $n_{\text{add}}$ while the dashed line represents the standard quantum limit of $0.5$ photons for a quantum limited amplifier.
We also measured the saturation power of the JPA (also called the $1$ dB compression point), which is the power of the input signal tone at which the gain decreases by $1$ dB. To carry out this measurement, we varied the power of the input signal tone while keeping the external magnetic flux, pump power, and pump frequency at the optimal values as mentioned above. The measured saturation power was around $-114$ dBm over the bandwidth of the JPA, as shown in Fig. 2(c).
@@
The equivalent circuit diagram of our device is shown in Fig. 3, where the impedance engineering is achieved by placing a series LC circuit (Transformer) between the transmission line (TL) and the JPA. In this diagram, $E_{J}$ and $C_{J}$ represent the Josephson energy and capacitance of the JPA, while $ME_{T}$ and $C_{T}$ correspond to the Josephson energy of each junction in the $M$-junction chain and capacitance of the transformer. The Hamiltonian for the device is
@@
$\hat{H}_{J}=-\frac{\hbar\Omega_{J}}{2}(\hat{A}_{J}-{\hat{A}_{J}}^{\dagger})^{2}-E_{J}\cos(c_{2}(\hat{A}_{J}+{\hat{A}_{J}}^{\dagger})),$ (3)
@@
$\hat{H}_{T}$ $=-\frac{\hbar\Omega_{T,\text{eff}}}{2}(\hat{A}_{T}-{\hat{A}_{T}}^{\dagger})^{2}$
@@
[…]
@@
[…]
@@
[…]
@@
Using our theory model, we have numerically computed the gain profile as follows. First, we solve Eq. 10 and Eq. 11 for the steady state solution starting with $\alpha_{T}(t=0)=\alpha_{J}(t=0)=0$ with a specified input pump power, $\alpha_{\text{in}}$ and a set of device parameters to obtain the long-time limit solutions $\alpha_{T}(t\rightarrow\infty),\alpha_{J}(t\rightarrow\infty)$. These steady-state pump occupations are then substituted in Eq. 12 and Eq. 13, which we use to calculate the gain. The device parameters used in the simulations are $\omega_{p}/{2\pi}=5.347\text{ GHz}$, $\Omega_{J}/{2\pi}=6.5\text{ GHz}$, $\Omega_{T,\text{eff}}/{2\pi}=6.218\text{ GHz}$, $L_{J}=0.37\text{ nH}$, and $L_{T}=1.9\text{ nH}$, where the relevant inductances are the sums of linearized inductance due to the junctions and geometric inductance due to the device geometry. The extra capacitance due to coupling to the environment was calculated for the experimental setup with a 444\,4GHz bandwidth, external impedance $R\sim 50\,\Omega$, and $C_{T}\sim 420\,\text{fF}$. In addition, the input pump power was set to align with the experimental conditions.
@@
In summary, we have fabricated and characterized a broadband impedance-engineered JPA in a single-step fabrication procedure. We measured a gain of 181818\,18dB over a 400 MHz bandwidth with near quantum-limited noise performance and saturation power of -114dBm. Our experimental results are comparable to state-of-the-art values reported in literature - impedance engineering with gain 202020\,20dB, bandwidth 640640640\,640MHz and saturation power $-110\,$dBm Roy et al. (2015) and impedance matching with gain 151515\,15dB, bandwidth 700700700\,700MHz and saturation power $-108\,$dBm Mutus et al. (2014) - while using a much simpler device fabrication process. The simpler fabrication process enables the faster and more reliable realisation of near-quantum-limited broadband amplifiers for multi-qubit readout Vijay (2020) or other sensing applications. Further extensions to this work can involve improving the saturation power by using an array of SQUIDs in the JPA and increasing the number of junctions in the transformer array to decrease transformer non-linearity. Furthermore, similar devices can be designed which use three-wave (instead of four-wave) mixing, which due to the large frequency separation of the pump and signal allow the pump tone to be easily filtered without affecting the amplified signal. Such devices can hence be used as a broadband squeezed microwave source for high-precision measurements where noise in one quadrature is below vacuum noiseMehmet et al. (2010). The compact footprint of the devices makes its on-chip integration into a qubit architecture possible, thereby mitigating insertion loss due to chip-to-chip interconnects Eddins et al. (2019).
@@
  • Hatridge et al. (2013) M. Hatridge, S. Shankar, M. Mirrahimi, F. Schackert, K. Geerlings, T. Brecht, K. M. Sliwa, B. Abdo, L. Frunzio, S. M. Girvin, R. J. Schoelkopf,  and M. H. Devoret, “Quantum back-action of an individual variable-strength measurement,” Science 339, 178–181 (2013)https://www.science.org/doi/pdf/10.1126/science.1226897 .
@@
  • Feldman, Parrish, and Chiao (1975) M. J. Feldman, P. T. Parrish,  and R. Y. Chiao, “Parametric amplification by unbiased Josephson junctions,” Journal of Applied Physics 46, 4031–4042 (1975)https://pubs.aip.org/aip/jap/article-pdf/46/9/4031/7955853/4031_1_online.pdf .
  • Wahlsten, Rudner, and Claeson (1977) S. Wahlsten, S. Rudner,  and T. Claeson, “Parametric amplification in arrays of Josephson tunnel junctions,” Applied Physics Letters 30, 298–300 (1977)https://pubs.aip.org/aip/apl/article-pdf/30/6/298/7738589/298_1_online.pdf .
@@
See pages 1, of Figures/supplementary.pdf See pages 2, of Figures/supplementary.pdf See pages 3, of Figures/supplementary.pdf See pages 4, of Figures/supplementary.pdf See pages 5, of Figures/supplementary.pdf See pages 6, of Figures/supplementary.pdf See pages 7, of Figures/supplementary.pdf See pages 8, of Figures/supplementary.pdf See pages 9, of Figures/supplementary.pdf See pages 10, of Figures/supplementary.pdf See pages 11, of Figures/supplementary.pdf

### v2

@@
We present the experimental demonstration of an impedance-engineered Josephson parametric amplifier (IEJPA) fabricated in a single-step lithography process. Impedance engineering is implemented using a lumped-element series LC circuit. We use a simpler lithography process where the entire device – impedance transformer and JPA – are patterned in a single electron beam lithography step, followed by a double-angle Dolan bridge technique for Al-AlOx-Al deposition. We observe amplification with 18 dB gain over a wide 400400\,MHz bandwidth centered around 5.35.3\,GHz with added noise approaching the quantum limit, and a saturation power of 114-114\,dBm. To accurately explain our experimental results, we extend existing theories for impedance-engineered JPAs to incorporate the full sine nonlinearity of both the JPA and the transformer. Our work shows a path to simpler realization of broadband JPAs and provides a theoretical foundation for a regime of JPA operation that has been less explored in literature.
@@
Josephson Parametric Amplifiers (JPAs) Yurke et al. (1989) have been a mainstay of superconducting circuit QED systems Vijay, Slichter, and Siddiqi (2011); Hatridge et al. (2013); Lin et al. (2013); Roy and Devoret (2016) due to their ability to provide quantum-limited amplificationCaves (1982), facilitate single-shot readoutKrantz et al. (2016), generate squeezed states of lightMallet et al. (2011); Yurke et al. (1988), and to generate two-mode squeezed states through three-wave and four-wave mixing mechanismsEichler et al. (2011). The phase-insensitive amplification mode of these amplifiers is usually characterized by their gain-bandwidth product defined by the maximum gain times the full width at half maximum of the amplifier. Conventional JPAs are usually narrow-band amplifiers with low gain-bandwidth products (2020\,dB gain over 205020-50\,MHz) Castellanos-Beltran and Lehnert (2007); Yamamoto et al. (2008); Mutus et al. (2013); Zhou et al. (2014), which limit their utility to amplifying a signal composed of a single mode/tone (frequency component). On the other hand, amplifiers with higher gain-bandwidth products can be used to amplify a range of frequencies over a wider bandwidth. This finds applications in multiplexing qubit readout by allowing the amplification of multiple readout tones within the broad bandwidth of the JPA Vijay (2020).
@@
In order to increase the bandwidth without compromising on the gain, and thereby boost the gain-bandwidth product, Travelling Wave Parametric Amplifiers (TWPAs) CULLEN (1958); Feldman, Parrish, and Chiao (1975); Wahlsten, Rudner, and Claeson (1977); Ho Eom et al. (2012); Yaakobi et al. (2013); Zorin (2019) were developed. TWPAs are transmission lines with a distributed nonlinear inductance that typically provide amplification of around 2020\,dB over a 44\,GHz bandwidth. However, they involve complex fabrication techniques that are required to create thousands of identical unit cells with Josephson junctions (JJs) and capacitors Macklin et al. (2015); White et al. (2015).
@@
In this Letter, we demonstrate an impedance-engineered JPA optimized for four-wave mixing-based amplification that is fabricated using a single-step lithography technique. The simplicity of our process, in turn, reduces the cost and time needed to fabricate the device. It also reduces the errors in fabrication and thereby increases the reliability and yield of the devices. The device used in this work has a simple design made up entirely of coplanar elements. Operating the device in reflection mode, we demonstrate a high gain over a wide bandwidth with noise performance approaching the quantum limit, comparable to state-of-the-art devices Roy et al. (2015); Grebel et al. (2021); Lu et al. (2022). Furthermore, to explain our results, we develop a theoretical model to compute the expected gain of an impedance-engineered JPA accounting for the full non-linearity of both the JPA and the impedance transformer. Our model is a substantial extension of previous theoretical work where only the lowest order non-linearity of the JPA is considered and the transformer is assumed to be made of a linear inductor.
@@
The device was fabricated on an intrinsic silicon substrate using an aluminum evaporative deposition process. First, we did a preliminary cleaning of the substrate using “piranha” (a $3:1$ volume-ratio mixture of $\text{H}_{2}\text{SO}_{4}$ and $\text{H}_{2}\text{O}_{2}$, for 10 minutes) followed by dilute hydrofluoric acid(HF) – a $50:1$ volume-ratio mixture of DI water and HF, for 30 seconds. Next, a resist bi-layer comprising a 500500\,nm thick bottom layer of methacrylic acid (MAA) in methyl methacrylate (MMA), diluted in ethyl lactate (EL) at a 11% concentration, and a 200200\,nm thick top layer of 4% PMMA 950K in Anisole was spin-coated on the substrate. We then carried out e-beam lithography in a Raith e-Line at 2020\,kV acceleration voltage. We pattern both the impedance transformer and the JPA in this step. The pattern development was done using the MIBK (Methyl Isobutyl Ketone) developer. Subsequently, we used the Dolan-bridge technique for forming the Josephson junctions through the double-angle (shadow) evaporation of aluminum. The double-angle deposition was done at $\pm 45^{\circ}$ with the first and second layers of Al having thicknesses 2020\,nm and 4040\,nm, respectively. All the Josephson junctions in the circuit – both in the JPA SQUID loop and the impedance transformer – were formed in a single in-situ oxidation step between the double-angle Al depositions. The oxidation was done at 700 mTorr pressure for 20 minutes.
The device was packaged inside a copper sample box using gold wire bonds. The sample box was then mounted on the mixing chamber plate of a dilution refrigerator at 1515\,mK. To tune the frequency of the JPA, we applied a DC magnetic flux using an external coil made of superconducting wire. The sample box containing the JPA device was mounted at the center of the external coil such that the plane of the SQUID of the JPA was perpendicular to the direction of the magnetic field of the coil. The resonance frequency $\Omega_{J}$ of the JPA could then be varied using the external magnetic flux $\Phi_{\text{DC}}$. To further prevent noise due to external magnetic fields, the sample box, along with the magnetic coil, was placed inside a superconducting aluminum shield mounted at the mixing chamber.
We measured the gain and the added noise of the JPA by sweeping the frequency of the signal tone in steps of 1.5 MHz over a broad frequency range ($\sim\,$1 GHz) and perfoming pump on/ pump off measurements using a spectrum analyser (details of the measurement setup can be found in the Supplementary material). We optimized the gain-bandwidth product of the JPA by tuning the external DC magnetic flux, pump frequency and pump power. We obtained an optimal maximum gain of 1818\,dB over a 400400\,MHz bandwidth, at a pump frequency of 5.3475.347\,GHz, and a pump power of 88-88\,dBm, as shown in Fig. 2(a). The ripple in the gain profile can be attributed to standing waves in the cable connecting the JPA input to the slightly impedance mismatched circulator port. The ripple can be reduced by using a shorter cable between the JPA and the circulator Roy et al. (2015).
@@
$n_{\text{add}}=LT_{\text{HEMT}}\Bigl[\frac{\text{SNR}_{\text{off}}}{\text{SNR}_{\text{on}}}-\frac{1}{G}\Bigr]\frac{k_{b}}{\hbar\omega},$ (1)
where $n_{\text{add}},T_{\text{HEMT}}$ are the added noise photons by JPA and linearised HEMT noise temperature respectively, $\text{SNR}_{\text{off}},\text{SNR}_{\text{on}}$ are the measured SNR when the pump is off and on respectively, $L$ is the loss between the JPA output and the HEMT input, and $G$ is the gain of the JPA. In our analysis, we assume that the $T_{\text{HEMT}}$ is equal to the nominal noise temperature in the datasheet of 3.63.6\,K. We also estimate the loss between the IEJPA and the HEMT to be about 1.81.8\,dB, mostly attributed to insertion loss of the circulators and the loss from the cables. The plot of estimated $n_{\text{add}}$ is shown by the orange curve in the middle panel of Fig. 2. We note that for a more precise characterization of the added noise, we need to use a shot-noise tunnel junction (SNTJ) noise thermometerSimoen et al. (2015), which was not available during this work. The green dotted line in Fig. 2(b) corresponds to the standard quantum limit of 0.50.5\, photons of noise added.
Figura 2: Measured (a) gain, (b) added noise, and (c) 1dB compression point for the IEJPA operated at a pump frequency and pump power of 5.3425.342\,GHz and 88-88\,dBm respectively. In the plot (b), the solid line represents the estimated $n_{\text{add}}$ while the dashed line represents the standard quantum limit of $0.5$ photons for a quantum limited amplifier.
We also measured the saturation power of the JPA (also called the $1$ dB compression point), which is the power of the input signal tone at which the gain decreases by $1$ dB. To carry out this measurement, we varied the power of the input signal tone while keeping the external magnetic flux, pump power, and pump frequency at the optimal values as mentioned above. The measured saturation power was around 114-114 dBm over the bandwidth of the JPA, as shown in Fig. 2(c).
@@
The equivalent circuit diagram of our device is shown in Fig. 3, where the impedance engineering is achieved by placing a series LC circuit (Transformer) between the transmission line (TL) and the JPA. In this diagram, $E_{J}$ and $C_{J}$ represent the Josephson energy and capacitance of the JPA, while $ME_{T}$ and $C_{T}$ correspond to the Josephson energy of each junction in the $M$-junction chain and capacitance of the transformer. In considering this effective circuit model we make some physically motivated simplifying assumptions which takes into account the effect of higher order nonlinearities of the device without making the device circuit too complex. We assume that for our device, the contribution from geometric inductance can be considered by just modifying the Josephson energies, or equivalently the linearized inductance, of the junctions. We justify this by noting that the geometric inductance contribution is about $15\%$ of the total linearized inductance for both the transformer and the JPA at the operating point. Another assumption that we make is that the direct capacitance from the transformer capacitance pads to ground can be neglected. This is justified from finite-element simulations that we performed that provide a capacitance to ground from the pads to be about $10\%$ of the effective capacitance between the capacitor pads, which only leads to a small frequency shift that can be accounted for by $C_{T}$. The Hamiltonian for the device is
@@
$\hat{H}_{J}=-\frac{\hbar\Omega_{J}}{4}(\hat{A}_{J}-{\hat{A}_{J}}^{\dagger})^{2}-E_{J}\cos(c_{2}(\hat{A}_{J}+{\hat{A}_{J}}^{\dagger})),$ (3)
@@
$\hat{H}_{T}$ $=-\frac{\hbar\Omega_{T,\text{eff}}}{4}(\hat{A}_{T}-{\hat{A}_{T}}^{\dagger})^{2}$
@@
$\hat{H}_{\text{int}}$ $=\hbar\int_{0}^{\infty}\sqrt{\frac{\omega R}{2\pi Z_{T}}}(\hat{A}_{T}(\omega)-{\hat{A}_{T}}^{\dagger}(\omega))(\hat{b}(\omega)-{\hat{b}^{\dagger}(\omega))}\,d\omega,$ (5)
@@
$\hat{H}_{\text{int}}=\hbar\int_{0}^{\infty}\sqrt{\frac{\kappa}{2\pi}}(\hat{A}_{T}(\omega)-{\hat{A}_{T}}^{\dagger}(\omega))(\hat{b}(\omega)-{\hat{b}^{\dagger}(\omega))}\,d\omega,$ (6)
@@
$\hat{H}_{\text{env}}=\hbar\int_{0}^{\infty}\omega{\hat{b}^{\dagger}(\omega)}\hat{b}(\omega)\,d\omega,$ (7)
@@
Using our theory model, we have numerically computed the gain profile as follows. First, we solve Eq. 10 and Eq. 11 for the steady state solution starting with $\alpha_{T}(t=0)=\alpha_{J}(t=0)=0$ with a specified input pump power, $\alpha_{\text{in}}$ and a set of device parameters to obtain the long-time limit solutions $\alpha_{T}(t\rightarrow\infty),\alpha_{J}(t\rightarrow\infty)$. These steady-state pump occupations are then substituted in Eq. 12 and Eq. 13, which we use to calculate the gain. The device parameters used in the simulations are $\omega_{p}/{2\pi}=5.347\text{ GHz}$, $\Omega_{J}/{2\pi}=6.5\text{ GHz}$, $\Omega_{T,\text{eff}}/{2\pi}=6.218\text{ GHz}$, $L_{J}=0.37\text{ nH}$, and $L_{T}=1.9\text{ nH}$, where the relevant inductances are the sums of linearized inductance due to the junctions and geometric inductance due to the device geometry. The extra capacitance due to coupling to the environment was calculated for the experimental setup with a 44\,GHz bandwidth, external impedance $R\sim 50\,\Omega$, and $C_{T}\sim 420\,\text{fF}$. In addition, the input pump power was set to align with the experimental conditions.
@@
In summary, we have fabricated and characterized a broadband impedance-engineered JPA in a single-step fabrication procedure. We measured a gain of 1818\,dB over a 400 MHz bandwidth with noise approaching the quantum limit and saturation power of -114dBm. Our experimental results are comparable to state-of-the-art values reported in literature - impedance engineering with gain 2020\,dB, bandwidth 640640\,MHz and saturation power 110-110\,dBm Roy et al. (2015) and impedance matching with gain 1515\,dB, bandwidth 700700\,MHz and saturation power 108-108\,dBm Mutus et al. (2014) - while using a much simpler device fabrication process. The simpler fabrication process enables the faster and more reliable realisation of near-quantum-limited broadband amplifiers for multi-qubit readout Vijay (2020) or other sensing applications. Further extensions to this work can involve improving the saturation power by using an array of SQUIDs in the JPA and increasing the number of junctions in the transformer array to decrease transformer non-linearity. Furthermore, similar devices can be designed which use three-wave (instead of four-wave) mixing, which due to the large frequency separation of the pump and signal allow the pump tone to be easily filtered without affecting the amplified signal. Such devices can hence be used as a broadband squeezed microwave source for high-precision measurements where noise in one quadrature is below vacuum noiseMehmet et al. (2010). The compact footprint of the devices makes its on-chip integration into a qubit architecture possible, thereby mitigating insertion loss due to chip-to-chip interconnects Eddins et al. (2019).
Supplementary Material
See the supplementary material for the details of the measurement setup for the gain and the added noise of the IEJPA. Also, we present a detailed description of our theory which includes the derivation of full Hamiltonian of the IEJPA, followed by pump and signal equations for Transformer and JPA.
@@
  • Hatridge et al. (2013) M. Hatridge, S. Shankar, M. Mirrahimi, F. Schackert, K. Geerlings, T. Brecht, K. M. Sliwa, B. Abdo, L. Frunzio, S. M. Girvin, R. J. Schoelkopf, and M. H. Devoret, “Quantum back-action of an individual variable-strength measurement,” Science 339, 178–181 (2013), https://www.science.org/doi/pdf/10.1126/science.1226897 .
@@
  • Feldman, Parrish, and Chiao (1975) M. J. Feldman, P. T. Parrish, and R. Y. Chiao, “Parametric amplification by unbiased josephson junctions,” Journal of Applied Physics 46, 4031–4042 (1975), https://pubs.aip.org/aip/jap/article-pdf/46/9/4031/7955853/4031_1_online.pdf .
  • Wahlsten, Rudner, and Claeson (1977) S. Wahlsten, S. Rudner, and T. Claeson, “Parametric amplification in arrays of josephson tunnel junctions,” Applied Physics Letters 30, 298–300 (1977), https://pubs.aip.org/aip/apl/article-pdf/30/6/298/7738589/298_1_online.pdf .